# Stable Quantum Vortices in Lee-Huang-Yang Dipolar Superfluids

### v1

@@ I Introduction
The experimental observation of dipolar Bose-Einstein condensates (BECs) in ultracold gases of 52Cr [3, 2, 1], 164Dy [5, 4] and 168Er [6] magnetic atoms has been a giant step forward in the studies of quantum matter [7, 8]. BECs of magnetic atoms exhibit long-range dipole-dipole interaction (DDI) between atoms, depending on the mutual orientation of atomic magnetic moments [9], in addition to nonlinear effects induced by the $s$-wave contact interactions between the atoms, including repulsion and attraction. On the other hand, techniques based on the Feshbach resonance can efficiently control the contact nonlinearity [10], enabling the control of the interplay between the nonlocal (DDI) and local (contact) interactions for performing various manipulations with the dipolar BECs. In particular, this setting was elaborated in Ref. [11] to control spatial separation of spin-coupled BECs. In this regard, it is relevant to note that tunable long-range interatomic correlations can be induced by laser-induced DDIs in BECs [12], creating a roton minimum in the excitation spectrum, reminiscent of the well-known properties of the strongly correlated superfluid represented by liquid helium II [13, 14]. Rotons are also well known as elementary excitations in dipolar BECs, see Refs. and references therein [15]-[18].
@@
In terms of the corresponding GP formalism, the correction to the MF approximation, as considered in Ref. [28], is represented by quartic self-repulsion term, which is usually much weaker than the cubic MF one. The LHY correction becomes significant in the special case when the intraspecies self-repulsion and interspecies cross-attraction terms in the corresponding system of nonlinearly coupled GP equations (mutuality fitted by means of the Feshbach resonance [35]) almost cancel each other for the symmetric state, with equal wave functions of the two components [28]. In particular, the “LHY superfluid” was predicted [36] and observed [37], in which the full cancellation of the MF fields makes the entire nonlinearity in the GP system represented solely by the LHY terms. In agreement with the predictions, the LHY corrections have made it possible to create robust QDs in binary BECs with the contact interactions [38, 39, 40, 41, 42], as well as in dipolar condensates dominated by the long-range magnetic DDIs (see theoretical predictions in Refs. [43, 44] and experimental realizations in Refs. [45, 46, 47]). In the latter case, the interplay between long-range interactions and the LHY effect also makes it possible to realize supersolids in the dipolar quantum gases [48, 49, 50], see also review [51]. The effects of quantum fluctuations may support supersolids in other settings as well [52, 53, 54, 55, 56, 57]. The role of the LHY corrections has also been demonstrated in experimental studies of the critical BEC temperature [58], quantum depletion  [59], excitation spectrum [60], and thermodynamic equation of state [61]. It is relevant to mention that, when investigating the universality of the contact interactions in the regime of observable beyond-MF effects, no measurable contributions from three-body interactions were found [62].
@@
Quantized vortices, first predicted in the framework of GP equations for trapped Bose gases in Ref. [63], have been created in experiments using rotating magnetic traps or applying laser stirring to BEC above a certain critical rotational frequency. The observation of the vortices provides compelling evidence for the occurrence of superfluidity in BECs, sharing some features with liquid helium [64]. In dipolar Bose gases, quantized vortices have been intensively investigated, within the MF approximation, in the last decades up to now [65, 66, 70, 67, 68, 69, 71, 72, 73, 74, 75, 80, 81, 76, 77, 78, 79]. The structure and stability of vortices in dipolar BECs are strongly affected by the anisotropy of the DDIs and their strengths, with respect to the contact interactions [65, 70, 71, 82]. Static hydrodynamic solutions and their stability have been also explored for the rotating dipolar BECs in the Thomas-Fermi (TF) limit [83]. In this context, the structure of dipolar BECs was studied in the quasi-2D scenario with arbitrary orientations of the dipoles [84, 73]. The second-order-like phase transition of straight and helical vortex lines, occurring due to the effect of the dipoles’ orientation, has also been reported [68]. DDI strongly affects the number, structure, and stability of vortices. Further, the critical rotational frequency for the nucleation of vortices decreases with the increase of the DDI strength, while the contact interactions enhance the vortex stability [74] . The stabilization of 2D and 3D self-trapped (localized) QDs with embedded vorticity due to the action of the LHY effect in the binary BECs with contact interactions were predicted, respectively, in Refs. [85] and [86]. On the other hand, the QD solutions for axially symmetric vortex states in the single-component dipolar BEC were found to be fully unstable [87].
@@ II The GP system with the DDI and LHY terms
$a_{\mathrm{dd}}\equiv m\mu_{0}\mu_{M}^{2}/(12\pi\hbar^{2}),$ (5)
@@ II.1 DDI in the quasi-2D BEC
$+$ $g_{\mathrm{dd}}(\varphi)\int\frac{d^{2}\mathbf{k}_{\rho}}{(2\pi)^{2}}e^{-i\mathbf{k}_{\rho}.\boldsymbol{\rho}}\widetilde{n}(\mathbf{k}_{\rho},t)h_{\mathrm{2D}}\biggr(\frac{k_{\rho}}{\sqrt{2\lambda}}\biggr)\psi,$
where $g_{\mathrm{dd}}(\varphi)\equiv(g_{\mathrm{dd}}/2){(3\cos^{2}\!\varphi-1)}$. Here, the coefficient $\eta$, which is defined below by Eq. (15), corresponds to the reduced LHY dimensionless coefficient $g_{\mathrm{LHY}}$ given by Eq. (4) for the dipolar system. The other factors in Eq. (II.1) are [95]
@@
$+$ $\frac{g_{\mathrm{dd}}(\varphi)}{2}\int\frac{d^{2}\mathbf{k}_{\rho}}{(2\pi)^{2}}e^{-i\mathbf{k}_{\rho}.\boldsymbol{\rho}}\widetilde{n}(\mathbf{k}_{\rho},t)h_{\mathrm{2D}}\biggr(\frac{k_{\rho}}{\sqrt{2\lambda}}\biggr)|\psi|^{2}\Bigg\}.$
@@ II.3 The pure LHY nonlinearity in the quasi-2D GP equation
This possibility to effectively cancel the local and nonlocal MF nonlinear terms are somewhat similarity to that proposed in Ref. [36], where a binary Bose mixture without DDI and rotation was considered, therefore our studies of on rotational properties of LHY superfluid, following below in section III.3, may be accordingly applied to the setting addressed in Ref. [36]. In addition, it is relevant to mention that, for the 2D BEC trapped in the harmonic-oscillator potential, the critical rotational frequency $\Omega_{c}$ for a single vortex production has been analytically determined in Ref. [100], assuming the strong cubic nonlinearity obeying the Thomas-Fermi limit. In the present notation, the corresponding expression is
@@ III Numerical results and discussion
The main objective of this work is to explore the nucleation and dynamics of vortices in the BEC dominated by the LHY nonlinearity, assuming the effective mutual cancellation of the MF contact interactions and DDIs. In subsection A, the results are produced in the absence of the LHY term, aiming to verify the rotation effect in the usual situation, where the MF nonlinearity is taken into account. These results, which are consistent with previous investigations, serve as a benchmark for the comparison with the cases where the LHY term is the dominant one. In that subsection, the explicit dependence on the orientation angle $\varphi$ is dropped in the DDI and LHY parameters, assuming that $\eta\equiv\eta(\varphi)$ is given by Eq. (16) and ${g}_{\mathrm{dd}}\equiv g_{\mathrm{dd}}(\varphi)$. Unless explicitly stated otherwise, this implies $\varphi=0^{\circ}$. In particular, the pure LHY superfluid considered in subsection C relies on the assumption of the effective cancellation of the MF contact and DDI terms, making it relevant to keep the $\varphi$ dependence of the $\eta$ and $g_{\mathrm{dd}}$ coefficients.
@@ III.1 Vortex modes in the absence of the LHY term
The numerical procedure to solve the 2D GP equation starts with the split-step Crank-Nicholson method. Its advantage is that the non-derivative nonlinear and linear terms can be treated very accurately, to secure the precision and robustness of the numerical scheme. In the present study, we have used the numerical code with stepsizes $\Delta x=$ $\Delta y=0.04$ and $\Delta t=0.001$ [94, 96, 101].
@@
The numerical results are exemplified by the vortex formation in two cases when the MF nonlinear terms in the GP equation represent the contact and dipolar interactions. The obtained results are consistent with previously reported ones, such as those for the rotating systems with the pure contact nonlinearity [63], and for the 3D dipolar rotating system with a high value of the aspect ratio [102]. Here, it is not only illustrative consider these results, which are also useful to estimate effects produced by the MF terms in comparison to those produced by the LHY nonlinearity. For this purpose, we keep the same DDI and contact-interaction parameters throughout the present study. Figure 2 shows how the vortices emerge following the variation of the rotation frequency $\Omega$ in the usual nonlinear MF system, with $g=100$, in the absence of the DDI and LHY terms ($g_{\mathrm{dd}}=\eta=0$) in Eq. (II.1). Consistent with the critical frequency (18) for the single vortex in the 2D BEC trapped in the harmonic-oscillator potential with frequency $\omega_{\perp}$ (where $\mu\approx 2.58$ is the scaled chemical potential), it is observed that the generation of vortices starts slightly below $\Omega\approx 0.6$, with the number of vortices $N_{v}$ increasing almost linearly in the interval of $0.7<\Omega<0.9$. In Fig 3, our focus is on the generation of vortices with the nonlinearity represented solely by the DDI term, so that $(g,g_{\mathrm{dd}},\eta)=(0,100,0)$. Comparing the results displayed in Figs. 2 and  3, one can understand the efficiency of the respective parameters for the control of the production of vortices. These two cases, presented in Figs. 2 and 3, in the absence of the LHY nonlinearity, will be useful for comparison with the impact of the LHY term on the vortex production, which is the subject of the next subsection.
Figure 3: (Color online) Density profiles $|\psi|^{2}$ projected onto the 2D plane, displaying vortex lattices with the nonlinearity represented solely by DDI with $g_{\mathrm{dd}}=100$. The rotation frequencies $\Omega$ are indicated in the panels, with the density levels coded by the color bars. The spatial domain and units are the same as in Fig. 2.
@@ III.2 Vorticity in dipolar BEC with contact and LHY nonlinearities
The effect of LHY correction to the MF GP equation is estimated in this section by assuming $\eta=200$ in Eq.(II.1), with $g=100$ and without dipolar interactions. By strictly considering the expression (15), we should be aware that any specific value assumed for $\eta$ will depend on $a_{s}$ and $a_{\mathrm{dd}}$. In particular, when $a_{s}=0$, the quantum fluctuations become zero, such that $\eta\neq 0$ with $a_{s}=0$ has no interest in our present analyses. However, we can verify rotational effects due to quantum fluctuation in a non-dipolar BEC system. For such a case, by assuming $g_{\mathrm{dd}}=0$ in Eq.  (II.1), we first consider the results obtained in Fig. 2 with $g=100$ to estimate how strong the parameter $\eta$ of the quantum fluctuations could be assumed to reach some visible effect in the results. By looking for that, we have fixed $\eta=200$ in our following studies, with the results being presented in Fig. 4. The quantum fluctuation contribution to the results, as expected, is quite small when compared with the MF nonlinear one, such that it starts to be noticeable only with $\eta=200$ at $\Omega$ between 0.5 and 0.6. This can be verified in our results of Figs. 2 and  4. The effect of quantum fluctuations with such $\eta$ can also be noticed for large rotations, close to $\Omega=0.99$.
@@ III.3 The generation of vortices in the pure-LHY superfluid
The rotating dipolar LHY superfluid is exemplified by setting $\eta=200$ in Eq. (II.3). For this dipolar BEC system, the values of $a_{s}$ and $a_{\mathrm{dd}}$, which are included in expression (15), are not critical, provided that they are tuned to secure the approximate cancellation of the local and nonlocal MF terms. By changing the rotation frequency $\Omega$ from 0, by increments of $\Delta\Omega=0.1$, we have observed no vortex produced for $\Omega\leq 0.6$ (in units of $\omega_{\perp}$), with a pair of vortices being verified at $\Omega=0.7$. Next, two and four pairs ($N_{v}=4$ and $8$) are observed at $\Omega=0.8$ and $\Omega=0.9$, respectively. These results are shown in Fig. 6. In more detail, we explored the values of $\Omega$ to produce the first four vortices, with results for the density plots and corresponding phases presented in Fig. 7. In this regard, it was verified that, to produce one or three stable vortices, careful fine tuning of $\Omega$ is required, with the values of $\Omega$ being, respectively, $0.6401$ and $0.768$. These findings reveal criticality in the states concerning $\Omega$, for relatively small number of produced vortices, $N_{v}$. More explicitly, at $\Omega=0.6400$, no vortex can be found, whereas two vortices are generated at $0.6410\leq\Omega\leq 0.767$, as shown in Fig. 7(b). Next, by slightly increasing the rotation frequency to $\Omega=0.768$, a state with three-vortex is obtained, which persists in a very narrow interval, $0.768\leq\Omega 0.778$. A small variation in $\Omega$ can transform the three-vortex state into a two- or four-vortex one (at $\Omega\leq 0.767$ and $\Omega\geq 0.779$, respectively. The four-vortex state is produced in a broader interval, $0.779\leq\Omega\leq 0.881$, with Fig. 8(d) displaying the result for $\Omega=0.779$. This behavior, with vortex configurations being more stable with $2$ and $4$ vortices than with $1$ and $3$ vortices, is apparently a characteristics of the pure LHY superfluid in our quasi-2D approach, which is represented by the quartic nonlinearity in the GP equation with a strong pancake-like aspect ratio.
Figure 7: (Color online) For the LHY superfluid with $\eta=200$ at four values of $\Omega$ (in units $\omega_{\perp}$), the density profiles $|\psi|^{2}$, with respective phases patterns shown in two upper rows, with the corresponding 3D representations drawn in the lower rows part of the figure. Panel (a) pertains to the critical frequency, $\Omega_{c}=0.6401$, for the creation of the single-vortex configuration. Coordinates $(x,y)$ cover intervals [5,+5][-5,+5] (in units of $\ell_{\perp}$).
@@
Figure 9 summarizes the results for $N_{v}$ versus $\Omega$, varying $\Omega$ up to $0.99$, for the solely-LHY nonlinearity, with strength $\eta=200$. In this case, the largest number of vortices found for $\Omega=0.99$ is $N_{v}=20$. The figure displays the range intervals of frequencies $\Delta\Omega$ for the given specific values of $N_{v}$. Notable are the cases with $N_{v}=2,4$ and $8$, which have larger $\Delta\Omega$ than the other cases. Also, noticiable is the case with $N_{v}=1$, which has $\Omega_{c}=0.6401$ with $\Delta\Omega<0.001$ (In this regard, see also the upper row of Fig.7). Inside Fig. 9, we have also included density plots obtained for $N_{v}=5,6,7$ and $10$, to complement the ones previously shown for $N_{v}\leq 4$.

### v2

@@ I Introduction
The experimental observation of dipolar Bose-Einstein condensates (BECs) in ultracold gases of 52Cr [3, 2, 1], 164Dy [5, 4] and 168Er [6] magnetic atoms has been a giant step forward in the studies of quantum matter [7, 8]. BECs of magnetic atoms exhibit long-range dipole-dipole interaction (DDI) between atoms, depending on the mutual orientation of atomic magnetic moments [9], in addition to nonlinear effects induced by the $s$-wave contact interactions between the atoms, including repulsion and attraction. On the other hand, techniques based on the Feshbach resonance can efficiently control the contact nonlinearity [10], enabling the control of the interplay between the nonlocal (DDI) and local (contact) interactions for performing various manipulations with the dipolar BECs. In particular, this setting was elaborated in Ref. [11] to control spatial separation of spin-coupled BECs. In this regard, it is relevant to note that tunable long-range interatomic correlations can be induced by laser-induced DDIs in BECs [12], creating a roton minimum in the excitation spectrum, reminiscent of the well-known properties of the strongly correlated superfluid represented by liquid helium II [13, 14]. Rotons are also well known as elementary excitations in dipolar BECs, see [15]-[18] and references therein.
@@
In terms of the corresponding GP formalism, the correction to the MF approximation, as considered in Ref. [28], is represented by quartic self-repulsion term, which is usually much weaker than the cubic MF one. The LHY correction becomes significant in the special case when the intraspecies self-repulsion and interspecies cross-attraction terms in the corresponding system of non-linearly coupled GP equations (mutuality fitted by means of the Feshbach resonance [35]) almost cancel each other for the symmetric state, with equal wave functions of the two components [28]. In particular, the “LHY superfluid” was predicted [36] and observed [37], in which the full cancellation of the MF fields makes the entire nonlinearity in the GP system represented solely by the LHY terms. In agreement with the predictions, the LHY corrections have made it possible to create robust QDs in binary BECs with the contact interactions [38, 39, 40, 41, 42], as well as in dipolar condensates dominated by the long-range magnetic DDIs (see theoretical predictions in Refs. [43, 44] and experimental realizations in Refs. [45, 46, 47]). In the latter case, the interplay between long-range interactions and the LHY effect also makes it possible to realize supersolids in the dipolar quantum gases [48, 49, 50], see also review [51]. The effects of quantum fluctuations may support supersolids in other settings as well [52, 53, 54, 55, 56, 57]. The role of the LHY corrections has also been demonstrated in experimental studies of the critical BEC temperature [58], quantum depletion  [59], excitation spectrum [60], and thermodynamic equation of state [61]. It is relevant to mention that, when investigating the universality of the contact interactions in the regime of observable beyond-MF effects, no measurable contributions from three-body interactions were found [62].
@@
Quantized vortices, first predicted in the framework of GP equations for trapped Bose gases in Ref. [63], have been created in experiments using rotating magnetic traps or applying laser stirring to BEC above a certain critical rotational frequency. The observation of the vortices provides compelling evidence for the occurrence of superfluidity in BECs, sharing some features with liquid helium [64]. In dipolar Bose gases, quantized vortices have been intensively investigated, within the MF approximation, in the last decades up to now [65, 66, 70, 67, 68, 69, 71, 72, 73, 74, 75, 80, 81, 76, 77, 78, 79]. The structure and stability of vortices in dipolar BECs are strongly affected by the anisotropy of the DDIs and their strengths, with respect to the contact interactions [65, 70, 71, 82, 83]. Static hydrodynamic solutions and their stability have been also explored for the rotating dipolar BECs in the Thomas-Fermi (TF) limit [84]. In this context, the structure of dipolar BECs was studied in the quasi-2D scenario with arbitrary orientations of the dipoles [85, 73]. The second-order-like phase transition of straight and helical vortex lines, occurring due to the effect of the dipoles’ orientation, has also been reported [68]. DDI strongly affects the number, structure, and stability of vortices. Further, the critical rotational frequency for the nucleation of vortices decreases with the increase of the DDI strength, while the contact interactions enhance the vortex stability [74] . The stabilization of 2D and 3D self-trapped (localized) QDs with embedded vorticity due to the action of the LHY effect in the binary BECs with contact interactions were predicted, respectively, in Refs. [86] and [87]. On the other hand, the QD solutions for axially symmetric vortex states in the single-component dipolar BEC were found to be fully unstable [88].
@@ II The GP system with the DDI and LHY terms
$a_{\mathrm{dd}}\equiv\frac{m\mu_{0}\mu_{M}^{2}}{12\pi\hbar^{2}},$ (5)
@@ II.1 DDI in the quasi-2D BEC
$+$ $g_{dd}(\varphi)\int\frac{d^{2}\mathbf{k}_{\rho}}{(2\pi)^{2}}e^{-i\mathbf{k}_{\rho}.\boldsymbol{\rho}}\widetilde{n}(\mathbf{k}_{\rho},t)h_{\mathrm{2D}}\biggr(\frac{k_{\rho}}{\sqrt{2\lambda}}\biggr)\psi,$
where $g_{dd}(\varphi)\equiv(g_{dd}/2){(3\cos^{2}\!\varphi-1)}$. Here, the coefficient $\eta$, which is defined below by Eq. (15), corresponds to the reduced LHY dimensionless coefficient $g_{\mathrm{LHY}}$ given by Eq. (4) for the dipolar system. The other factors in Eq. (II.1) are [96]
@@
$+$ $\frac{g_{dd}(\varphi)}{2}\int\frac{d^{2}\mathbf{k}_{\rho}}{(2\pi)^{2}}e^{-i\mathbf{k}_{\rho}.\boldsymbol{\rho}}\widetilde{n}(\mathbf{k}_{\rho},t)h_{\mathrm{2D}}\biggr(\frac{k_{\rho}}{\sqrt{2\lambda}}\biggr)|\psi|^{2}\Bigg\}.$
@@ II.3 The pure LHY nonlinearity in the quasi-2D GP equation
This possibility to effectively cancel the local and nonlocal MF nonlinear terms is somewhat similarity to that proposed in Ref. [36], where a binary Bose mixture without DDI and rotation was considered, therefore our studies of on rotational properties of LHY superfluid, following below in section III.3, may be accordingly applied to the setting addressed in Ref. [36]. In addition, it is relevant to mention that, for the 2D BEC trapped in the harmonic-oscillator potential, the critical rotational frequency $\Omega_{c}$ for a single vortex production has been analytically determined in Ref. [101], assuming the strong cubic nonlinearity obeying the Thomas-Fermi limit. In the present notation, the corresponding expression is
@@ III Numerical results and discussion
The main objective of this work is to explore the nucleation and dynamics of vortices in the BEC dominated by the LHY nonlinearity, assuming the effective mutual cancellation of the MF contact interactions and DDIs. In subsection A, the results are produced in the absence of the LHY term, aiming to verify the rotation effect in the usual situation, where the MF nonlinearity is taken into account. These results, which are consistent with previous investigations, serve as a benchmark for the comparison with the cases where the LHY term is the dominant one. In that subsection, the explicit dependence on the orientation angle $\varphi$ is dropped in the DDI and LHY parameters, assuming that $\eta\equiv\eta(\varphi)$ is given by Eq. (16) and ${g}_{\mathrm{dd}}\equiv g_{\mathrm{dd}}(\varphi)$. Unless explicitly stated otherwise, this implies $\varphi=0^{\circ}$. In particular, the pure LHY superfluid considered in subsection C relies on the assumption of the effective cancellation of the MF contact and DDI terms, making it relevant to keep the $\varphi$ dependence of the coefficients $\eta$ and $g_{dd}$.
@@ III.1 Vortex modes in the absence of the LHY term
The numerical procedure to solve the 2D GP equation starts with the split-step Crank-Nicholson method. Its advantage is that the non-derivative nonlinear and linear terms can be treated very accurately, to secure the precision and robustness of the numerical scheme. In the present case, the usual split-step Crank-Nicholson method was combined with the fast Fourier transform approach, following the prescription given in Refs. [83, 95, 97]. Here, we have used a numerical code with 512 $\times$ 512 grid size, with spatial step sizes $\Delta x=$ $\Delta y=0.04$. The norm of the wave function is restored to be 1 at each time step $\Delta t=0.01$.
@@
The numerical results are exemplified by the vortex formation in two cases when the MF nonlinear terms in the GP equation represent the contact and dipolar interactions. The obtained results are consistent with previously reported ones, such as those for the rotating systems with the pure contact nonlinearity [63], and for the 3D dipolar rotating system with a high value of the aspect ratio [102]. Here, it is not only illustrative to consider these results, which are also useful to estimate effects produced by the MF terms in comparison to those produced by the LHY nonlinearity. For this purpose, we keep the same DDI and contact-interaction parameters throughout the present study. Figure 2 shows how the vortices emerge following the variation of the rotation frequency $\Omega$ in the usual nonlinear MF system, with $g=100$, in the absence of the DDI and LHY terms ($g_{dd}=\eta=0$) in Eq. (II.1). Consistent with the critical frequency (18) for the single vortex in the 2D BEC trapped in the harmonic-oscillator potential with frequency $\omega_{\perp}$ (where $\mu\approx 2.58$ is the scaled chemical potential), it is observed that the generation of vortices starts slightly below $\Omega\approx 0.6$, with the number of vortices $N_{v}$ increasing almost linearly in the interval of $0.7<\Omega<0.9$. In Fig 3, our focus is on the generation of vortices with the nonlinearity represented solely by the DDI term, so that $(g,g_{dd},\eta)=(0,100,0)$. Comparing the results displayed in Figs. 2 and  3, one can understand the efficiency of the respective parameters for the control of the production of vortices. These two cases, presented in Figs. 2 and 3, in the absence of the LHY nonlinearity, will be useful for comparison with the impact of the LHY term on the vortex production, which is the subject of the next subsection.
Figure 3: (Color online) Density profiles $|\psi|^{2}$ projected onto the 2D plane, displaying vortex lattices with the nonlinearity represented solely by DDI with $g_{dd}=100$. The rotation frequencies $\Omega$ are indicated in the panels, with the density levels coded by the color bars. The spatial domain and units are the same as in Fig. 2.
@@ III.2 Vorticity in dipolar BEC with contact and LHY nonlinearities
The effect of LHY correction to the MF GP equation is estimated in this section by assuming $\eta=200$ in Eq.(II.1), with $g=100$ and without dipolar interactions. By strictly considering the expression (15), we should be aware that any specific value assumed for $\eta$ will depend on $a_{s}$ and $a_{\mathrm{dd}}$. In particular, when $a_{s}=0$, the quantum fluctuations become zero, such that $\eta\neq 0$ with $a_{s}=0$ has no interest in our present analyses. However, we can verify rotational effects due to quantum fluctuation in a non-dipolar BEC system. For such a case, by assuming $g_{dd}=0$ in Eq.  (II.1), we first consider the results obtained in Fig. 2 with $g=100$ to estimate how strong the parameter $\eta$ of the quantum fluctuations could be assumed to reach some visible effect in the results. By looking for that, we have fixed $\eta=200$ in our following studies, with the results being presented in Fig. 4. The quantum fluctuation contribution to the results, as expected, is quite small when compared with the MF nonlinear one, such that it starts to be noticeable only with $\eta=200$ at $\Omega$ between 0.5 and 0.6. This can be verified in our results of Figs. 2 and  4. The effect of quantum fluctuations with such $\eta$ can also be noticed for large rotations, close to $\Omega=0.99$.
@@ III.3 The generation of vortices in the pure-LHY superfluid
The rotating dipolar LHY superfluid is exemplified by setting $\eta=200$ in Eq. (II.3). For this dipolar BEC system, the values of $a_{s}$ and $a_{\mathrm{dd}}$, which are included in expression (15), are not critical, provided that they are tuned to secure the approximate cancellation of the local and nonlocal MF terms. By changing the rotation frequency $\Omega$ from 0, by increments of $\Delta\Omega=0.1$, we have observed no vortex produced for $\Omega\leq 0.6$ (in units of $\omega_{\perp}$), with a pair of vortices being verified at $\Omega=0.7$. Next, two and four pairs ($N_{v}=4$ and $8$) are observed at $\Omega=0.8$ and $\Omega=0.9$, respectively. These results are shown in Fig. 6. In more detail, we explored the values of $\Omega$ to produce the first four vortices, with results for the density plots and corresponding phases presented in Fig. 7. The phase map uses a color scale from $\pi$ to $\pi$, determined by the $\arctan({\rm Im}\psi/{\rm Re}\psi)$. Beyond the trap boundaries, the phase is a numerical artifact, which arises from densities that are numerically small but not vanish.
As observed by the $\Omega$ values indicated in the four panels of Fig. 7, particularly in the cases to produce one and three stable vortices, careful fine tuning of $\Omega$ was required, with the values of $\Omega$ being, respectively, $0.6401$ and $0.768$. These findings reveal criticality in the states concerning $\Omega$, for relatively small number of produced vortices, $N_{v}$. More explicitly, at $\Omega=0.6400$, no vortex can be found, whereas two vortices are generated at $0.6410\leq\Omega\leq 0.767$, as shown in Fig. 7(b). Next, by slightly increasing the rotation frequency to $\Omega=0.768$, a state with three-vortex is obtained, which persists in a very narrow interval, $0.768\leq 0.778$. A small variation in $\Omega$ can transform the three-vortex state into a two- or four-vortex one (at $\Omega\leq 0.767$ and $\Omega\geq 0.779$, respectively. The four-vortex state is produced in a broader interval, $0.779\leq\Omega\leq 0.881$, with Fig. 8(d) displaying the result for $\Omega=0.779$. This behavior, with vortex configurations being more stable with $2$ and $4$ vortices than with $1$ and $3$ vortices, is apparently a characteristics of the pure LHY superfluid in our quasi-2D approach, which is represented by the quartic nonlinearity in the GP equation with a strong pancake-like aspect ratio.
Figure 7: (Color online) For the LHY superfluid with $\eta=200$ at four values of $\Omega$ (in units $\omega_{\perp}$), the density profiles $|\psi|^{2}$ shown in the first row with the respective phases patterns shown in second row while the corresponding 3D representations drawn in the lower rows of the figure. Panel (a) pertains to the critical frequency, $\Omega_{c}=0.6401$, for the creation of the single-vortex configuration. The phase color map (the second row) spans from $\pi$ (darkest blue) to $\pi$ (brightest yellow), encircling each singularity in a clockwise direction, with zero value denoted by the maximum-red color.
@@
Figure 9 summarizes the results for $N_{v}$ versus $\Omega$, varying $\Omega$ up to $0.99$, for the solely-LHY nonlinearity, with strength $\eta=200$. In this case, the largest number of vortices found for $\Omega=0.99$ is $N_{v}=20$. The figure displays the range intervals of frequencies $\Delta\Omega$ for the given specific values of $N_{v}$. Notable are the cases with $N_{v}=2,4$ and $8$, which have larger $\Delta\Omega$ than the other cases. Also, noticeable is the case with $N_{v}=1$, which has $\Omega_{c}=0.6401$ with $\Delta\Omega<0.001$ (In this regard, see also the upper row of Fig.7). Inside Fig. 9, we have also included density plots obtained for $N_{v}=5,6,7$ and $10$, to complement the ones previously shown for $N_{v}\leq 4$.